OpenStudy (anonymous):

A coat is on sale for Rs.120 after a discount of 20 percent. Find the original price.

OpenStudy (anonymous):

do u know how to find 20% of 120 ??

OpenStudy (kirbykirby):

Assume the original price to be \(x\). If you discount 20 percent of \(x\), you are essentially subtracting off \(0.20x\) from \(x\). So you can formulate this as: \(x - 0.20x = 120 \\ 0.80x = 120 \\ x = 120/0.80\\ x = 150\) You can double check: 20% of 150 is (150 * 0.20 = 30) And 150 - 30 = 120 :)